# Numerical aspects of integration in physics-informed neural networks modelling

### v1

@@ Abstract
In this paper we numerically analyse problems that arise in numerical integration used in various problems that are nowadays being solved by physics-informed neural networks. In particular we show how inaccurately evaluated integrands can lead to severe precision problems. The tricky and serious part of the problem is especially in its detection, when for example a simple macroscopic visualization of the integrand does not indicate any potential problems and obtained results are often wrongly treated as correct. To solve the problem, it is crucial to decide if it is sufficient to evaluate the integrand in the standard double arithmetic or if a higher precision arithmetic has to be used. We compare and recommend numerical quadratures for selected examples and different parameter values, especially for problematic cases.
Received 9 August 2024

### v2

# Challenges in numerical integration in physics-informed neural networks modelling

@@ Abstract
In this paper, we numerically examine the challenges that emerge in numerical integration for various tasks now tackled by physics-informed neural networks. Specifically, we illustrate how inaccurately computed integrands can cause serious precision issues. A major difficulty lies in detecting these problems, since a simple large-scale view of the integrand may not reveal any potential errors, and the resulting outcomes are often mistakenly considered correct. To address this, it is critical to determine whether standard double-precision arithmetic suffices for evaluating the integrand or if higher precision is necessary. We evaluate and suggest numerical quadrature methods for selected examples and parameter values, particularly focusing on challenging scenarios.
Received 6 February 2025
@@ 6 Conclusion
Numerical integration challenges, as discussed in this paper, arise due to inaccuracies in evaluating integrands, leading to severe precision problems in many models, and as demonstrated in particular in physics-informed neural networks. These issues stem from floating-point arithmetic limitations and improper numerical quadrature methods, especially when model parameters exacerbate these problems. Similar challenges can occur for example in the peridynamic models, which rely on integral formulations to capture long-range interactions and avoid spatial derivatives. The evaluation of integro-differential operators in peridynamics, as seen in its bond-based models, requires precise numerical integration for accurate representation of material behavior, including crack propagation and heterogeneous material response (Haghighat, Bekar, Madenci, and Juanes, 2021; Ning, Cai, Dong, Liu, and Wang, 2023).
In PINNs applied to peridynamics, numerical integration problems may emerge when enforcing physical constraints, such as integro-differential operators, within the loss function. For example, in peridynamic-informed PINNs, inaccurate integration of forces across horizons or improperly handled kernel functions can degrade model accuracy, particularly under complex boundary or material conditions (Ning, Cai, Dong, Liu, and Wang, 2023; Difonzo, Lopez, and Pellegrino, 2024). These problems echo those highlighted in this paper, where detection of precision loss can be non-trivial without detailed numerical diagnostics. Thus, addressing these integration challenges in peridynamic models often necessitates adopting adaptive quadrature methods or higher-precision arithmetic to maintain robustness and accuracy (Ning, Cai, Dong, Liu, and Wang, 2023).